OpenStudy (anonymous):

PLEASE HELP! Write the quadratic equation given the solutions x=-3 and x=4. How would I do this, step by step please?

OpenStudy (anonymous):

If the roots are x = -4 and x = 3, then the equation would be: (x+4) * (x-3) = 0 Multiply these factors together to get: x^2 + x - 12 = 0 Since your c factor has to be 4, then you need to divide both sides of this equation by -3 to get: -x^2/3 - x/3 + 4 = 0 This would be the same as: -(1/3)*x^2 - (1/3)*x + 4 = 0 a would be equal to -(1/3) b would be equal to -(1/3) c would be equal to 4 If you were given this equation in this form to start with and asked to solve for the roots, you would have done the following: Multiply both sides of the equation by (-3) to get: x^2 + x - 12 = 0 Factor this equation to get: (x-3) * (x+4) = 0 Solve for x to get: x = 3 or: x = -4 which is where you started from.

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

No problem! :)